# Online Energy Storage Arbitrage under Imperfect Predictions: A Conformal Risk-Aware Approach

### v1

@@ 3.4. Problem Statement
A flowchart of the proposed approach is summarized in Figure 1. First, as in Remark 2, we establish that imperfect predictions inevitably lead to suboptimal decisions and potential profit losses for energy storage arbitrage. Though such profit losses are inherently unobservable in real-time operations, Proposition 1 demonstrates that they can be estimated using the temporal difference error as a proxy. Therefore, the key insight underlying our proposed approach is to constrain the arbitrage cumulative risk, quantified through the observable temporal difference error (referred to as the loss of interest in the following sections). Controlling this cumulative risk then serves as an indirect mechanism for managing the unobservable profit risk. This risk management approach is achieved through a conformal decision-making framework that dynamically calibrates a conformal control variable, with detailed methodology presented in Section 4 and the main theoretical results in Theorem 2. The online calibration framework supports various practical loss metrics implementable in real-time decision-making, such as temporal prediction error, temporal value error, etc. A comprehensive discussion of these loss functions and their practical implications is provided in Section 5.
@@ 4. Conformal Risk-Aware Energy Storage Arbitrage
The arbitrage problem (1) describes a risk-neutral framework. We now consider an arbitrage strategy in a risk-aware context that mitigates downside risk, utilizing the conformal decision theory (Lekeufack et al., 2024). In this section, we first briefly summarize the main results of this theoretical framework and then establish its compatibility in the context of our problem.
@@ 4.1. Conformal Decision Theory
The dispatch decision $D_{t}^{\gamma_{t}}$ is parameterized by $\gamma_{t}$, which controls the level of conservativeness with higher values of $\gamma_{t}$ corresponding to more aggressive strategies (as specified later in Section 4.2). We evaluate the quality of the decision-making through a loss function $\mathcal{L}(\cdot)$ [^2] The framework works for any bounded loss; however, to simplify analysis, we assume the loss is normalized to the interval [0, 1].. Typically, the potential loss increases when aggressive decisions are taken (large $\gamma_{t}$), while overly small choices of $\gamma_{t}$ leads to conservative and under-performing decisions. This trade-off often results in monotone or near-monotone loss functions with respect to $\gamma$.
@@
Conformal controllers (Lekeufack et al., 2024) offer a simple yet effective way to address this challenge for storage arbitrage. They calibrate dispatch decisions by adjusting the control parameter $\gamma_{t}$ to maintain statistical guarantees on realized losses of interest under uncertainty. Let $\mathcal{T}^{+}$ denote the set of all time steps $t\in\{1,\ldots,T+1\}$. Building on conformal controller design from (Lekeufack et al., 2024, Theorem 1), we extend its context to energy storage arbitrage and present the following conformal controller. The proof is also adapted accordingly and revised for clarity in Appendix C.
@@
Theorem 2 establishes the boundedness of cumulative risk for energy storage online arbitrage. It is important to emphasize that this bounded risk does not directly correspond to profit loss, but rather serves as a measurable and effective proxy for it, as formalized in Proposition 1. We’ll explore this relationship further in the following sections.
@@
Though the conformal controller (10) defined in Theorem 2 doesn’t explicitly require any monotonicity in the storage arbitrage loss function $\ell_{t}$, this strategy becomes practically effective only if the loss is monotone or near-monotone in the conformal control variable $\gamma_{t}$. Otherwise, the assumption on eventual safety in Definition 1 will not hold. More details on conformal risk control and the monotonicity of loss functions can be found in (Angelopoulos et al., 2023a). The monotonicity of the loss functions in our storage arbitrage algorithm will be discussed in more detail in Section 5, where we introduce two choices of loss function.
@@ 4.2. Risk-Aware Arbitrage using Conformal Decision Theory
where $\sigma$ is a fixed sensitivity parameter, $z_{\gamma}$ represents the quantile of level $\gamma$ of the normal distribution. The choice of $\sigma$ directly correlates with market volatility, enabling flexibility across different market conditions. At this point it is worth some time to relate the prediction set $\hat{C}_{t}$ to the accuracy of the forecasting model $\mathcal{M}$. If $\mathcal{M}$ was perfectly accurate then $\hat{C}_{t}$ would simply be the singleton $\hat{q}$. Not only is perfect forecasting not possible, but so is our ability to accurately quantify our forecasting error. Assume that $\epsilon_{t}$ is the ground truth forecast accuracy, where $\epsilon_{t}=0$ denotes a perfect forecast, in this setting we would have $\hat{C}_{t}=\hat{q}_{t}$. As $\epsilon_{t}$ increases, i.e., the forecast quality decreases, then $\hat{C}_{t}$ will also increase. As we cannot know $\epsilon_{t}$ we instead parametrize the prediction set by a proxy, $\gamma_{t}$, that reflects our confidence in the interval. We make the relationship between $\epsilon_{t}$ and $\gamma_{t}$ precise later in Proposition 2 and provide an illustrative example in Section 4.3.
@@
The conformal dispatch policy $\{D_{t}^{\gamma_{t}}\}_{t\in\mathcal{T}}$ applies the conformal prediction set (12) to the marginal opportunity value function forecasts $\hat{q}_{t}$ for each time step $t$ in the optimal arbitrage problem (1). The risk-neutral dispatch policy derived in (Zheng et al., 2022) and described in Appendix (17) then becomes:
@@
This strategy makes the unit more selective about trading opportunities, reflecting a risk-aware preference. While trading frequency decreases, the expected profit per trade increases, resulting in more conservative operation but higher confidence in profitability.
@@
Remark 4.
The conformal strategy with smaller $\gamma$ creates wider prediction sets, making decisions less sensitive to the underlying value function forecast differences.
The proposed risk-aware arbitrage strategy using the conformal controller is summarized in Algorithm 1. An illustrative example regarding arbitrage performance is provided in the following subsection, where we compare the conformal risk-aware strategy with a risk-neutral strategy. Specially, the unit can adjust its risk posture based on specific preferences—focusing either on the upper or lower side of the price spread—by modifying the construction of the conformal prediction set (12).
@@ 5.1. Strategy Calibration via Prediction Error
Following the discussion in Section 4.3, it’s natural to assume that each level of (observed) forecast accuracy corresponds to an appropriate prediction set width as in (12) that allows the storage to optimally balance risk and opportunity.
Proposition 0.
The proof of Proposition 2 is provided in Appendix D. This proposition establishes the correlation between forecast accuracy and the optimal size of the prediction set.
@@ 5.2. Strategy Calibration via Value Error
Due to the space limit, we defer the proof of Proposition 5 to Appendix E. Proposition 5 indicates a direct relationship between the step loss $\ell_{t}$ and the conformal control variable $\gamma$, which is a desirable property for risk-aware decision-making.
@@ 6.1.1. Data Description
We evaluate arbitrage strategy performance using real-time market price data from the New York ISO (NYISO) for the New York City (NYC) zone during 2023 (New York ISO, 2025). Value function forecaster training and uncertainty quantification are conducted using historical price data from 2021 to 2022. The price data Prices is recorded at 5-minute intervals.
@@ 6.1.2. Parameter Settings and Metrics
Our simulation features a storage unit with power rating $P=0.5MW$ and capacity $E=1.0MWh$ that starts and ends at 50% SoC, with discharging and charging efficiency $\eta=0.9$. Regarding the value function forecaster $\mathcal{M}$, we evaluate algorithm performance under two forecasting scenarios using the model configuration from (Baker et al., 2023): a good-accuracy forecaster ($R^{2}=0.4$) and a poor-accuracy forecaster ($R^{2}=-0.4$). This setup enables assessment of the algorithm robustness across varying prediction quality levels as measured by $R^{2}$ coefficients. For CC-prediction err, we construct the conservativeness calibration function in the form of $f:\hat{\gamma}_{t}\leftarrow\bar{\gamma}\left(1-\exp(-k\Delta q)\right)$. We use the cumulative profit as the performance metric: $\sum_{\tau=1}^{t}\lambda_{\tau}(p^{*}_{\tau}-b^{*}_{\tau})-cp^{*}_{\tau}$, which measures the total profit gained over the time horizon $\{1,\ldots,t\}$.
@@ 6.2.1. Baseline Algorithms
We evaluate the two variants of our proposed conformal risk-aware algorithm against the baseline algorithms summarized in Table 1, detailed in Appendix F. For benchmarking purposes, we also include the offline optimal profit, a non-causal upper bound computed in hindsight using the groundtruth prices and thus opportunity value functions.
@@ 6.2.2. Arbitrage Performance Comparison
Figure 4. Arbitrage performance comparison across different algorithms under varying opportunity value function forecast accuracy (measured by $R^{2}$): Offline Optimal represents the optimum assuming perfect knowledge; Risk Neutral deploys (1); CVaR sets risk scaling factor $\mu=1$; Chance Constrained assumes $\lambda_{t}$ following normal distribution and sets threshold $\Gamma=0.6$; Robust Optimization estimates an ellipsoid uncertainty set and sets threshold $\Gamma=1.0$; Switching Cost sets the penalty factor $\zeta=400$; CC-prediction err sets conformal controller update rate $\rho=0.001$ and the mapping function with $\hat{\gamma}=3.0$ and $k=0.1$; both CC-prediction err and CC-value err set $\rho=0.001$.
Figure 4 presents a comprehensive comparison of arbitrage performance across different algorithms under varying opportunity value function forecast accuracy levels.
With good forecast accuracy ($R^{2}=0.4$) as shown in Figure 4(4(a)), the proposed methods remain competitive with baseline approaches while providing additional risk protection. The risk-neutral strategy performs well under accurate predictions, as expected, but the conformal strategies maintain similar performance levels while offering downside protection. Notably, CC-prediction err closely tracks the risk-neutral strategy under good forecasting conditions, showing that the method adaptively reduces conservativeness when forecast accuracy improves. Overall, the conformal strategies demonstrate competitive performance under good forecasting conditions.
When forecast accuracy is poor ($R^{2}=-0.4$) as shown in Figure 4(4(b)), both proposed conformal strategies (CC-prediction err and CC-value err) significantly outperform all baseline methods. The conformal approaches recover approximately 47% of the offline optimal profit while the risk-neutral baseline incurs substantial losses, demonstrating the effectiveness of conformal risk control in managing uncertainty when predictions are unreliable. The conformal strategies exhibit superior performance under poor forecasting conditions.
@@ 6.2.3. Daily Decision Analysis
Under reliable predictions as shown in Figure 5(5(a)), CC-prediction err closely follows the offline optimal and risk-neutral strategies, particularly during high-value trading opportunities around hours 18-22 when prices peak above \$30/MWh. The strategy confidently captures discharge opportunities during price spikes while charging during low-price periods. CC-value err shows slightly more conservative behavior but maintains similar trading patterns.
@@ 6.2.4. Impact of Risk Preference
It is also important to emphasize that the temporal loss definition $\ell_{t}$, used to calculate cumulative risk, is calibration-method specific. Therefore, the relationship between profit loss and $\hat{R}_{T}$ is not guaranteed to be strictly monotonic across methods. For example, in Figure 6(6(b)), CC-value err yielding lower $\hat{R}_{T}$ does not necessarily lead to lower profit loss.
@@ Appendix A Conformal Prediction
Conformal prediction (Angelopoulos and Bates, 2022) is a distribution-free uncertainty quantification paradigm. In contrast to methods reliant on specific distributional or model assumptions, conformal prediction provides explicit, non-asymptotic statistical guarantees applicable to any pretrained prediction model $\hat{f}$, such as a neural network.
@@
The coverage guarantee is valid for any choice of model or score function, detailed proof available in (Angelopoulos and Bates, 2022, Appendix D). The selection of an appropriate score function, however, is critical as it dictates the efficiency and utility (e.g., the average size or adaptivity) of the resulting prediction sets. In practice, typical choices include absolute residuals in regression tasks or one minus the softmax probability of the true class in classification tasks.

### v2

@@ 3.4. Problem Statement
3.5. Proposed Approach Overview
A flowchart of the proposed approach is summarized in Figure 1. First, as in Section 3.4, we establish that imperfect predictions inevitably lead to suboptimal decisions and potential profit losses for energy storage arbitrage. Though such profit losses are inherently unobservable in real-time operations, they can be estimated using the temporal difference error as a proxy as later demonstrated in Proposition 3. Therefore, the key insight underlying our proposed approach is to constrain the arbitrage cumulative risk, quantified through the observable temporal difference error (referred to as the loss of interest in the following sections). Controlling this cumulative risk then serves as an indirect mechanism for managing the unobservable profit risk. This risk management approach is achieved through a conformal decision-making framework that dynamically calibrates a conformal control variable, with detailed methodology presented in Section 4 and the main theoretical results in Theorem 2. The online calibration framework supports various practical loss metrics implementable in real-time decision-making, such as temporal prediction error, temporal value error, etc. A comprehensive discussion of these loss functions and their practical implications is provided in Section 5.
@@ 4. Conformal Risk-Aware Energy Storage Arbitrage
The arbitrage problem (1) describes a risk-neutral framework. We now consider an arbitrage strategy in a risk-aware context that mitigates downside risk, utilizing the conformal decision theory (lekeufack2024). In this section, we first briefly summarize the main results of this theoretical framework and then establish its compatibility in the context of our problem.
@@ 4.1. Conformal Decision Theory
The dispatch decision $D_{t}^{\gamma_{t}}$ is parameterized by $\gamma_{t}$, which controls the level of conservativeness with higher values of $\gamma_{t}$ corresponding to more aggressive strategies (as specified later in Section 4.2). We evaluate the quality of the decision-making through a loss function $\mathcal{L}(\cdot)$. [^2] The framework works for any bounded loss; however, to simplify analysis, we assume the loss is normalized to the interval [0, 1]. Typically, the potential loss increases when aggressive decisions are taken (large $\gamma_{t}$), while overly small choices of $\gamma_{t}$ leads to conservative and under-performing decisions. This trade-off often results in monotone or near-monotone loss functions with respect to $\gamma$.
@@
Conformal controllers (lekeufack2024) offer a simple yet effective way to address this challenge for storage arbitrage. They calibrate dispatch decisions by adjusting the control parameter $\gamma_{t}$ to maintain statistical guarantees on realized losses of interest under uncertainty. Let $\mathcal{T}^{+}$ denote the set of all time steps $t\in\{1,\ldots,T+1\}$. Building on conformal controller design from (lekeufack2024, Theorem 1), we extend its context to energy storage arbitrage and present the following conformal controller. The proof is also adapted accordingly and revised for clarity in Appendix D.
@@
Theorem 2 establishes the boundedness of cumulative risk for energy storage online arbitrage. It is important to emphasize that this bounded risk does not directly correspond to profit loss, but rather serves as a measurable and effective proxy for it, as formalized below.
Proposition 0.
The proof of Proposition 3 is provided in Appendix C. This result suggests that the temporal difference error serves as a measurable proxy for value error, which is the true but typically unobservable objective. Here, this means the arbitrage cumulative risk serves as a proxy for the underlying profit loss risk. We’ll explore this relationship further in the following sections.
@@
Though the conformal controller (11) defined in Theorem 2 doesn’t explicitly require any monotonicity in the storage arbitrage loss function $\ell_{t}$, this strategy becomes practically effective only if the loss is monotone or near-monotone in the conformal control variable $\gamma_{t}$. Otherwise, the assumption on eventual safety in Definition 1 will not hold. More details on conformal risk control and the monotonicity of loss functions can be found in (angelopoulos2023risk). The monotonicity of the loss functions in our storage arbitrage algorithm will be discussed in more detail in Section 5,where we introduce two choices of loss function.
@@ 4.2. Risk-Aware Arbitrage using Conformal Decision Theory
where $\sigma$ is a fixed sensitivity parameter, $z_{\gamma}$ represents the quantile of level $\gamma$ of the normal distribution. The choice of $\sigma$ directly correlates with market volatility, enabling flexibility across different market conditions. At this point it is worth some time to relate the prediction set $\hat{C}_{t}$ to the accuracy of the forecasting model $\mathcal{M}$. If $\mathcal{M}$ was perfectly accurate then $\hat{C}_{t}$ would simply be the singleton $\hat{q}_{t}$. Not only is perfect forecasting not possible, but so is our ability to accurately quantify our forecasting error. Assume that $\epsilon_{t}$ is the ground truth forecast accuracy, where $\epsilon_{t}=0$ denotes a perfect forecast, in this setting we would have $\hat{C}_{t}=\hat{q}_{t}$. As $\epsilon_{t}$ increases, i.e., the forecast quality decreases, then $\hat{C}_{t}$ will also increase. As we cannot know $\epsilon_{t}$, we instead parametrize the prediction set by a proxy, $\gamma_{t}$, that reflects our confidence in the interval. We make the relationship between $\epsilon_{t}$ and $\gamma_{t}$ precise later in Proposition 4 and provide an illustrative example in Section 4.3.
@@
The conformal dispatch policy $\{D_{t}^{\gamma_{t}}\}_{t\in\mathcal{T}}$ applies the conformal prediction set (13) to the marginal opportunity value function forecasts $\hat{q}_{t}$ for each time step $t$ in the optimal arbitrage problem (1). The risk-neutral dispatch policy derived in (zheng2022) and described in Appendix (18) then becomes:
@@
This strategy makes the unit more selective about trading opportunities, reflecting a risk-aware preference. For example, using a smaller $\gamma$ in the conformal strategy creates wider prediction sets, which reduces trading frequency. Though fewer trades are taken, the expected profit per trade increases, resulting in more conservative operation but higher confidence in profitability. As a result, decision-making becomes less sensitive to the underlying value function forecast differences.
@@
Proposition 0.
The proof of Proposition 4 is provided in Appendix E. This proposition establishes that each level of (observed) forecast accuracy $\epsilon_{t}$ corresponds to an appropriate conformal control variable $\gamma_{t}$, and thus an appropriate prediction set width, as in (13), allowing the storage to optimally balance risk and opportunity.
The proposed risk-aware arbitrage strategy using the conformal controller is summarized in Algorithm 1. An illustrative example regarding arbitrage performance is provided in the following subsection, where we compare the conformal risk-aware strategy with a risk-neutral strategy.
@@ 5.2. Strategy Calibration via Value Error
Due to the space limit, we defer the proof of Proposition 4 to Appendix F. Proposition 4 indicates a direct relationship between the step loss $\ell_{t}$ and the conformal control variable $\gamma$, which is a desirable property for risk-aware decision-making.
@@ 6. Case Study
This section presents case studies evaluating the proposed conformal risk-aware arbitrage strategy under realistic real-time market conditions. We validate its risk management effectiveness by comparing it with representative baselines and examining performance under varying risk preferences. Extended sensitivity analyses, including broader parameter variations, are provided in Appendix H for completeness.
@@ 6.1.1. Data Description
We evaluate arbitrage strategy performance using real-time market price data from the New York ISO (NYISO) for the New York City (NYC) zone during 2023 (NYISO_price). Value function forecaster training and uncertainty quantification are conducted using historical price data from 2021 to 2022. The price data is recorded at 5-minute intervals.
@@ 6.1.2. Parameter Settings and Metrics
Our simulation features a storage unit with power rating $P=0.5MW$ and capacity $E=1.0MWh$ that starts and ends at 50% SoC, with discharging and charging efficiency $\eta=0.9$. Regarding the value function forecaster $\mathcal{M}$, we evaluate algorithm performance under two forecasting scenarios using the model configuration from (baker2023transferable): a good-accuracy forecaster ($R^{2}=0.4$) and a poor-accuracy forecaster ($R^{2}=-0.4$). This setup enables assessment of the algorithm robustness across varying prediction quality levels as measured by $R^{2}$ coefficients. For CC–prediction err, we construct the conservativeness calibration function in the form of $f:\hat{\gamma}_{t}\leftarrow\bar{\gamma}\left(1-\exp(-k\Delta q)\right)$. We use the cumulative profit as the performance metric: $\sum_{\tau=1}^{t}\lambda_{\tau}(p^{*}_{\tau}-b^{*}_{\tau})-cp^{*}_{\tau}$, which measures the total profit gained over the time horizon $\{1,\ldots,t\}$.
@@ 6.2.1. Baseline Algorithms
We evaluate the two variants of our proposed conformal risk-aware algorithm against the baseline algorithms summarized in Table 1, detailed in Appendix G. For benchmarking purposes, we also include the offline optimal profit, a non-causal upper bound computed in hindsight using the groundtruth prices and thus opportunity value functions. The detailed configuration details for all baseline algorithms are provided in the caption of Figure 4.
@@ 6.2.2. Arbitrage Performance Comparison
Figure 4. Arbitrage performance comparison across different algorithms under varying opportunity value function forecast accuracy (measured by $R^{2}$): Offline Optimal represents the optimum assuming perfect knowledge; Risk Neutral deploys (1); CVaR sets risk scaling factor $\mu=1$; Chance Constrained assumes $\lambda_{t}$ following normal distribution and sets threshold $\Gamma=0.6$; Robust Optimization estimates an ellipsoid uncertainty set and sets threshold $\Gamma=1.0$; Switching Cost sets the penalty factor $\zeta=400$; CC–prediction err sets the mapping function with $\hat{\gamma}=3.0$ and $k=0.1$; both CC–prediction err and CC–value err set $\rho=0.001$ and $\sigma=10$.
Figure 4 presents a comprehensive comparison of arbitrage performance across different algorithms under varying opportunity value function forecast accuracy levels: a reasonably good forecast ($R^{2}=0.4$) for general demonstration, and a poor forecast ($R^{2}=-0.4$) for extreme stress testing.
With good forecast accuracy ($R^{2}=0.4$) as shown in Figure 4(4(a)), the proposed methods remain competitive with baseline approaches while providing additional risk protection. The risk-neutral strategy performs well under accurate predictions, as expected, but the conformal strategies maintain similar performance levels while offering downside protection. Notably, CC–prediction err closely tracks the risk-neutral strategy under good forecasting conditions, showing that the method adaptively reduces conservativeness when forecast accuracy improves. Overall, the conformal strategies demonstrate competitive performance under good forecasting conditions.
When forecast accuracy is poor ($R^{2}=-0.4$) as shown in Figure 4(4(b)), both proposed conformal strategies (CC–prediction err and CC–value err) significantly outperform all baseline methods. The conformal approaches recover approximately 47% of the offline optimal profit while the risk-neutral baseline incurs substantial losses, demonstrating the effectiveness of conformal risk-aware decision-making in managing uncertainty when predictions are unreliable. The conformal strategies exhibit superior performance under poor forecasting conditions.
@@ 6.2.3. Daily Decision Analysis
Under reliable predictions as shown in Figure 5(5(a)), CC–prediction err closely follows the offline optimal and risk-neutral strategies, particularly during high-value trading opportunities around hours 18-22 when prices peak above \$30/MWh. The strategy confidently captures discharge opportunities during price spikes while charging during low-price periods. CC–value err shows slightly more conservative behavior but maintains similar trading patterns.
@@ 6.2.4. Impact of Risk Preference
It is also important to emphasize that the temporal loss definition $\ell_{t}$, used to calculate cumulative risk, is calibration-method specific. Therefore, the relationship between profit loss and $\hat{R}_{T}$ is not guaranteed to be strictly monotonic across methods. For example, in Figure 6(6(b)), CC–value err yielding lower $\hat{R}_{T}$ does not necessarily lead to lower profit loss.
7. Discussion and Future Work
7.1. Linear Degradation Modeling
Our framework currently models battery degradation as a constant marginal operational cost $C$. This linear approximation preserves the convexity of the optimization problem, enabling the derivation of the computationally efficient analytical dispatch policies in Eq. (14). While this simplification is standard in literature (xu2020; baker2023transferable; yi2025; wu2025; yi2025decision), we recognize that realistic battery degradation is often nonlinear, exhibiting increased marginal costs at extreme states of charge or higher discharge rates. Adopting a nonlinear model would alter the marginal opportunity value function, leading to smoother continuous dispatch responses rather than the threshold-based behavior observed under linear costs.
However, the overall robustness of the proposed conformal framework remains intact. First, the theoretical guarantees for risk control (Theorem 2) and the validity of the temporal difference error as a proxy (Proposition 3) rely on the boundedness and suitable monotonicity of the calibration loss, not on the specific form of the operational cost function. Second, the core idea of our conformal risk-aware framework remains applicable, as its role is to adjust decision conservativeness in response to forecast quality. That said, omitting nonlinear degradation may introduce systematic bias in the value function that the controller must counteract by widening prediction sets.
Future work will extend the conformal risk controller to accommodate these nonlinear or numerical optimization settings, thereby allowing the framework to simultaneously manage market price uncertainty and complex physical aging processes.
7.2. Symmetric Uncertainty Assumptions
For simplicity, the prediction set $\hat{C}_{t}$ (13) is constructed using symmetric quantile bounds, which assume approximately symmetric forecast errors. This design choice facilitates implementation and analytical tractability. However, given that electricity prices often exhibit skewed or log-normal distributions, the conformal risk management framework can be readily extended to more informative asymmetric prediction sets by selecting distinct upper and lower quantile levels according to specific risk preferences, as illustrated for example in (pouplin2025). We highlight this as an important direction for future refinement of the controller design.
@@ Appendix A Conformal Prediction
Conformal prediction (angelopoulos2022) is a distribution-free uncertainty quantification paradigm. In contrast to methods reliant on specific distributional or model assumptions, conformal prediction provides explicit, non-asymptotic statistical guarantees applicable to any pretrained prediction model $\hat{f}$, such as a neural network.
@@
The coverage guarantee is valid for any choice of model or score function, detailed proof available in (angelopoulos2022, Appendix D). The selection of an appropriate score function, however, is critical as it dictates the efficiency and utility (e.g., the average size or adaptivity) of the resulting prediction sets. In practice, typical choices include absolute residuals in regression tasks or one minus the softmax probability of the true class in classification tasks.